OpenStudy (anonymous):

What is the length of the missing side of the right triangle shown below? right triangle with legs measuring 5 and 3 units

OpenStudy (anonymous):

\[a^2 + b^2 = c^2\] where a = 5 and b = 3. Then you get 25 + 9 = 34 = c^2 Take sqrt and get c = \[\sqrt{34}\]

OpenStudy (anonymous):

Thanks you a lot

OpenStudy (anonymous):

do you understand how i did it?

OpenStudy (anonymous):

yes i do